The demand of a product is normally distributed with mean 1,000 and standard deviation 400. The overage cost is $30 and underage cost is $50.

a.) What is the optimal order quantity?

b.) What is the expected profit?

c.)The manager decided to order just the mean demand. What is the expected profit under this policy?
Answer #1

a.

P(R<Q*) = Cu/ (Cu+Co) = 50/(50+30) = 0.625

NORMSINV(0.625) gives 0.318639

Q* - mean / st. dev = 0.318639

Q* = 0.318639* 400 + 1000

Q* = 1127.4556

Q* = 1127

Due to HomeworkLib guidelines, we are not supposed to answer multiple questions. Please post individually. Thank you.